# Measurement of spin correlation and entanglement in ATLAS and CMS

### v1

@@ 5 Observation of quantum entanglement by CMS
Based on $pp$ collisions data from 2016 at $\sqrt{s}=13\,\mathrm{TeV}$ with an integrated luminosity of $36.3\,\mathrm{fb}^{-1}$, CMS measured entanglement in the $t\bar{t}$ dilepton channel near threshold [24]. The region is defined by $345<m_{t\bar{t}}<400\,\mathrm{GeV}$ and $\beta_{z}({t\bar{t}})<0.9$, where $\beta_{z}({t\bar{t}})$ is the relative longitudinal velocity between the laboratory and $t\bar{t}$ reference frames. This selection on $\beta_{z}({t\bar{t}})$ enhances the fraction of gluon–gluon fusion events, in which the singlet configuration dominates. The observable $D$ is extracted at parton level via a binned profile likelihood fit to the $\cos\varphi$ distribution, incorporating, for the first time in an LHC measurement, simplified quasi-bound state effects modelled as a pseudoscalar resonance $\eta_{t}$ [37]. Figure 4 shows the fitted and predicted values of $D$ with and without the $\eta_{t}$ contribution. While entanglement is observed with a significance exceeding $5\sigma$ in both with and without $\eta_{t}$ scenarios, the data is better described when $\eta_{t}$ is included in the signal model. The largest impact on the fit arises from the 50% normalisation applied to $\eta_{t}$ in the fit. Despite the simplified modelling, this analysis represents a step towards accounting for dynamics at threshold. Ongoing theoretical studies aim to refine these descriptions [40, 34, 33], and the threshold region has become increasingly important in light of recent ATLAS [12] and CMS [25] results where an excess compatible with a quasi–bound-state formation at threshold has been observed in both experiments.
@@ 6 Measurements of polarisation, spin correlations, and entanglement using lepton+jets events
CMS also performed measurements of quantum entanglement, including for the first time the high-$m_{t\bar{t}}$ (boosted) regime. Using the Peres–Horodecki criterion, an observable analogous to the threshold witness is defined as $\tilde{D}=3\cdot\frac{\langle\cos\chi\rangle}{\alpha_{1}\alpha_{2}}$, where $\chi$ differs from $\varphi$ by a sign flip of the $\hat{n}-$axis component of one of the spin analysers. Entanglement is indicated by $\tilde{D}>1/3$. A complementary approach uses full $t\bar{t}$ spin-density matrix reconstruction to define an entanglement-sensitive quantity $\Delta_{E}=C_{nn}+|C_{rr}+C_{kk}|$, with $\Delta_{E}>1$ signalling entanglement. In the boosted phase space defined by $m_{t\bar{t}}>800\,\mathrm{GeV}$ and $|\cos(\theta)|<0.4$, the matrix-based observable provides greater sensitivity than $\tilde{D}$, as illustrated in Figure 6. Owing to the limited event yield in this challenging region, the statistical uncertainty remains comparable in size to the systematic uncertainty. Nevertheless, this measurement clearly demonstrates that entanglement can be probed in the boosted regime, marking an important step towards studying high-energy $t\bar{t}$ spin correlations.

### v2

@@ 5 Observation of quantum entanglement by CMS
Based on $pp$ collisions data from 2016 at $\sqrt{s}=13\,\mathrm{TeV}$ with an integrated luminosity of $36.3\,\mathrm{fb}^{-1}$, CMS measured entanglement in the $t\bar{t}$ dilepton channel near threshold [26]. The region is defined by $345<m_{t\bar{t}}<400\,\mathrm{GeV}$ and $\beta_{z}({t\bar{t}})<0.9$, where $\beta_{z}({t\bar{t}})$ is the relative longitudinal velocity between the laboratory and $t\bar{t}$ reference frames. This selection on $\beta_{z}({t\bar{t}})$ enhances the fraction of gluon–gluon fusion events [2], in which the singlet configuration dominates. The observable $D$ is extracted at parton level via a binned profile likelihood fit to the $\cos\varphi$ distribution, incorporating, for the first time in an LHC measurement, simplified $t\bar{t}$ quasi-bound state effects. Such quasi-bound effects were first predicted in the late 1980s [37, 36, 45], nearly a decade before the discovery of the top quark at the Tevatron in 1995 [30, 19]. In this measurement, these effects are modelled as a pseudoscalar resonance $\eta_{t}$ [42]. Figure 4 shows the fitted and predicted values of $D$ with and without the $\eta_{t}$ contribution. While entanglement is observed with a significance exceeding $5\sigma$ in both with and without $\eta_{t}$ scenarios, the data is better described when $\eta_{t}$ is included in the signal model. The largest impact on the fit arises from the 50% normalisation applied to $\eta_{t}$ in the fit. Despite the simplified modelling, this analysis represents a step towards accounting for dynamics at threshold. Ongoing theoretical studies aim to refine these descriptions [46, 39, 38], and the threshold region has become increasingly important in light of recent ATLAS [13] and CMS [27] results where an excess compatible with a quasi–bound-state formation at threshold has been observed in both experiments.
@@ 6 Measurements of polarisation, spin correlations, and entanglement using lepton+jets events
CMS also performed measurements of quantum entanglement, including for the first time the high-$m_{t\bar{t}}$ (boosted) regime. Using the Peres–Horodecki criterion, an observable analogous to the threshold witness is defined as $\tilde{D}=3\cdot\frac{\langle\cos\chi\rangle}{\alpha_{1}\alpha_{2}}$, where $\chi$ differs from $\varphi$ by a sign flip of the $\hat{n}-$axis component of one of the spin analysers [2]. Entanglement is indicated by $\tilde{D}>1/3$. A complementary approach uses full $t\bar{t}$ spin-density matrix reconstruction to define an entanglement-sensitive quantity $\Delta_{E}=C_{nn}+|C_{rr}+C_{kk}|$, with $\Delta_{E}>1$ signalling entanglement. In the boosted phase space defined by $m_{t\bar{t}}>800\,\mathrm{GeV}$ and $|\cos(\theta)|<0.4$, the matrix-based observable provides greater sensitivity than $\tilde{D}$, as illustrated in Figure 6. Owing to the limited event yield in this challenging region, the statistical uncertainty remains comparable in size to the systematic uncertainty. Nevertheless, this measurement clearly demonstrates that entanglement can be probed in the boosted regime, marking an important step towards studying high-energy $t\bar{t}$ spin correlations.